Menu A restaurant offers a dinner special that has 12 choices for entrées,
10 choices for side dishes, and 6 choices for dessert. For the special, you can
choose one entrée, two side dishes, and one dessert. How many different
meals are possible?

Answer by stanbon(75887)

Menu A restaurant offers a dinner special that has 12 choices for entrées,
10 choices for side dishes, and 6 choices for dessert. For the special, you can
choose one entrée, two side dishes, and one dessert. How many different
meals are possible?
------------------
Ans: 12*10C2*6 = 12*45*6 = 3240 different meals
------------
Cheers,
Stan H.
============